### v2

# Cleaning Up the Streets: Understanding Motivations, Mental Models, and Concerns of Users Flagging Social Media Posts

Abstract.
Social media platforms offer flagging, a technical feature that empowers users to report inappropriate posts or bad actors to reduce online harm. The deceptively simple flagging interfaces on nearly all major social media platforms disguise complex underlying interactions among users, algorithms, and moderators. Through semi-structured interviews with 25 social media users experienced in flagging inappropriate content, we examine end-users’ understanding of flagging procedures, explore the factors that motivate them to flag, and surface their cognitive and privacy concerns. Our findings uncover a lack of procedural transparency in flagging mechanisms that create gaps in users’ mental models. Regardless, users strongly believe that it is crucial for platforms to provide flagging options. We highlight how flags raise questions about distributing labor and responsibility between platforms and users for addressing online harm. We recommend innovations in the flagging design space that assist user comprehension, ensure privacy, and reduce cognitive burdens.
content moderation, flags, transparency
copyright: acmcopyrightjournalyear: 2018doi: XXXXXXX.XXXXXXXjournal: JACMjournalvolume: 37journalnumber: 4article: 111publicationmonth: 8price: 15.00isbn: 978-1-4503-XXXX-X/18/06ccs: Human-centered computing Empirical studies in collaborative and social computing
1. Introduction
End-users have several options to take action when encountering inappropriate content on social media platforms, from regular interactions built into a platform (e.g., unfollowing an account) to content moderation-specific interactions. Of these options, it is only through flagging that end-users can directly notify platform administrators. Thus, the flag serves as a powerful and empowering tool for user-driven moderation.
Flagging a post involves three temporally distinct stages from the flagger’s perspective: (1) before, (2) during, and (3) after flagging. Each step grows more complex upon closer inspection. How does a user decide if the content is inappropriate? What do they do when platform guidelines are unclear? How do users interact with flag submission interfaces? What makes a flag submission effective, ensuring users feel their efforts weren’t wasted? In asking these questions, we are led to Crawford and Gillespie’s description of flagging as a “a complex interplay between users and platforms, humans and algorithms, and the social norms and regulatory structures of social media”  (Crawford and Gillespie, 2016). It is with this in mind that we examine the perspectives of users through these three steps of the flagging processes, ultimately aiming to determine how we may better understand flagging mechanisms to improve them for users.
Though at first glance, flag implementations may signal platforms’ commitment to enacting democratic governance and fostering user communities, in practice, flags lack transparency and are often designed without adequately considering users’ needs. Most major platforms that host user-generated content, e.g., Instagram (Fig. 1), thank users for flagging content, tell them they will be notified when a decision is made, and provide them with additional safety options such as blocking, muting, and hiding the flagged content. However, only Facebook offers users a centralized interface where they may revisit past flags to check their review status (Fig. 2).
(a) A thank you message to users for flagging a post on Instagram.
(b) Additional options to block and hide content after flagging on Instagram.
Figure 1. Instagram’s flagging interface pages.
(a) A list of submitted reports, as listed in Facebook’s support inbox.
(b) A thank you message to users for flagging a post on Facebook.
Figure 2. Facebook’s flagging interface pages.
In this study, we analyze user experiences with flagging mechanisms on major social media platforms and ask three key research questions. First, given the limited information cues regarding flagging procedures in all stages, it is unclear how users interpret them and what shapes their understanding. Therefore, we ask:
RQ1: What are users’ mental models of how platforms implement flagging mechanisms, and how do they develop these models?
Next, we examine why users choose to flag content. Prior research on a gaming platform (Kou and Gui, 2021) shows that gamers often flag content for reasons that do not align with the platform’s notion of toxicity (e.g., flagging teammates’ poor game performance). We focus on flagging behaviors in social media platforms and the factors that drive users to flag by asking:
RQ2: What motivates users to engage (or not engage) in flagging on social media sites?
Submitting a flag often involves traversing a multi-step interface that asks users to fit their objections with a post or account under the platforms’ pre-defined categories of inappropriate conduct. We explore the cognitive and privacy-related challenges of submitting a flag by asking:
RQ3: What are users’ cognitive and security concerns about engaging in flagging?
To answer these questions, we conducted semi-structured interviews with 25 social media users who recently flagged content. We first asked participants about their experiences of flagging inappropriate content online. Next, we asked questions that probed their understanding of how flagging works. Finally, we examined participants’ perspectives on possible changes to flagging mechanisms, especially regarding their transparency and feedback.
Applying interpretive qualitative analysis (Merriam, 2002) to our interview data, we found that the lack of transparency in flagging creates substantial gaps in users’ mental models about how platforms evaluate flagged content. However, users developed strategies to leverage their interpretation of how flags work, using flags alongside other site mechanisms, such as blocking and private messaging, to address online harms. A belief in generalized reciprocity motivates users to flag inappropriate content—they expect a collective flagging effort would help keep online spaces clean. However, flagging is not without its costs. We surfaced reservations users have about the cognitive burdens of flagging and privacy concerns when discussing potential innovations to flagging mechanisms.
Our main contributions are:
  1. (1)
    Empirical data on user experiences and perspectives about flagging on social media platforms. We provide empirical data contributing to Crawford and Gillespie’s theoretical conceptualization of flagging  (Crawford and Gillespie, 2016). Our findings highlight the importance of flags on social media as a vital way for users to voice concerns within an opaque content moderation system. We explore diverse views of flags—as a right, obligation, exploitation, and a commons problem. We argue that platforms should promote digital literacy around flagging without imposing it as a duty.
  2. (2)
    Conceptualization of the three stages of flagging (before, during, and after). We provide a conceptual understanding of flagging from end users’ perspectives across these three temporal stages. We also highlight the key user concerns within each stage.
  3. (3)
    Design opportunities with recommendations. We provide empirically informed design recommendations and identify opportunities for innovation. We recommend incorporating seams (Eslami et al., 2016) into flagging interfaces to help users understand their operations and ensure transparency throughout the flagging process. Innovations to flagging mechanisms should address user needs while respecting privacy and cognitive burden. Additionally, we identify opportunities for platforms to collaborate with marginalized communities to address the flagging-related challenges they face.
2. Related Work
2.1. Content Moderation Affordances on Social Media
Most contemporary platforms hosting user-generated content offer a mechanism for users to provide feedback on the content they see that violates the platform’s guidelines, known as a “flag.” Platforms use flags to identify potentially inappropriate content, which human moderators or bots subsequently review  (Kou and Gui, 2021). If this review determines that the content violates the platform’s guidelines, the flagged user may face direct consequences, such as content removal and account suspension  (Jhaver et al., 2019a, 2021).
We contextualize flags as part of a collection of mechanisms that let users express their voices in content moderation systems. These mechanisms commonly involve expressing dissatisfaction about types of content (e.g., hate speech, violence, spam)  (Kumar et al., 2021; Banko et al., 2020; Scheuerman et al., 2021) or behaviors (e.g., harassment, bullying, scamming)  (Shaw, 2018). Flags are often accompanied by a variety of personal moderation tools, which have been defined as “tools that let users configure or customize some aspects of their moderation preferences on social media”  (Jhaver et al., 2023). Personal moderation tools include account-based tools that let users mute or block specific accounts (Geiger, 2016; Matias et al., 2015) and content-based tools that let users hide comments containing configured keywords  (Jhaver et al., 2022) or set up content sensitivity controls  (Jhaver and Zhang, 2023). We focus on flagging, one of the most conspicuous content moderation mechanisms available to general end-users. While personal moderation tools change only the configuring user’s news feed, flags enable site-wide content removal, provide higher granularity in letting users express their dissatisfaction, and prompt users to justify their actions. Appeal mechanisms  (Vaccaro et al., 2020, 2021), which let moderated users report their dissatisfaction with content moderation decisions and request that they be reversed, are often used to contest platform moderation decisions about flagged content or other content removed by administrators. Users may also utilize the above moderation mechanisms in combination with regular platform interactions that demonstrate positive or negative algorithmic feedback (e.g., like and dislike buttons) (Khan, 2017) or other social signals (e.g., unfollowing accounts)  (Sasahara et al., 2021).
Users with special roles (e.g., community moderators) usually have access to a wider range of moderation tools  (Hatmaker, 2021; Stockinger et al., 2023). For example, Reddit moderators may independently program Automod, a bot, to automatically identify and take action against content that infringes upon their community’s guidelines (Wright, 2022; Jhaver et al., 2019b). Similarly, Twitch streamers and YouTube creators rely on third-party applications for harmful content detection (Cai and Wohn, 2019; Jhaver et al., 2022). We do not focus on such specialized tools because this study is examining regular end-users who lack access to them. However, we draw from literature on accountability and division of labor involved in using such tools  (Jhaver et al., 2019b; Anaobi et al., 2023; Chandrasekharan et al., 2019) to examine how end users perceive the use of automation and human moderators to review flagged content.
2.2. User Perspectives on Content Moderation
Prior work asserted the importance of examining user perspectives of content moderation to inform platform leadership and policymakers. Scholars argue that these internet governance regimes are defining users’ roles and shaping our political and public discourse, so it is imperative to incorporate users’ perspectives in their design  (Riedl et al., 2021; Langvardt, 2017). Our research on users’ perceptions of flagging mechanisms highlights existing user needs to inform changes to how platforms design and administer these systems.
The discussion of user perspectives includes questions of who owns the responsibility for platform governance. Gillespie advocated for a collective civic responsibility to govern platforms  (Gillespie, 2018). However, it is not always clear how we may distinguish specific stakeholder responsibilities to apply this form of collective governance. For instance, users have indicated that they attribute responsibility for interventions and detecting inappropriate content to platforms (Riedl et al., 2021) or human moderators  (Naab et al., 2018) rather than themselves. However, users were found to prefer personally configurable moderation tools that regulate hate speech, violent content, and sexually explicit content instead of platform-enacted bans of those categories (Jhaver and Zhang, 2023). We build on this work by examining how end-users perceive their responsibilities to flag fitting in with how flagging responsibilities are divided.
Prior research has highlighted substantial concerns about the burdens of physical and emotional labor that moderation places on regular end-users  (Jhaver et al., 2023), volunteer moderators (Li et al., 2022a; Jhaver et al., 2019b; Wohn, 2019; Dosono and Semaan, 2019; Ruckenstein and Turunen, 2020), and commercial moderators  (Roberts, 2019; Steiger et al., 2021) and the tensions of how users’ volunteer efforts factor into platforms’ monetary gains  (Wohn, 2019; Li et al., 2022b). While the emotional burden of moderation on volunteer and hired human moderators has been examined in detail (Dosono and Semaan, 2019; Roberts, 2019), we find that scant prior literature has observed the additional burden on regular users of reporting content and thus aim to explore this aspect in our study.
As an extension of discussions regarding the responsibilities to flag harmful content and the burden it places on users, it is known that people of marginalized identities are more likely to experience harm on online platforms (Duggan, 2017). For example, Black and female-identifying users were shown to disproportionately experience an array of harms from unwanted behaviors on platforms (e.g., hate speech, doxxing)  (Musgrave et al., 2022). To address these challenges, HCI and CSCW researchers have prioritized the content moderation needs and perspectives of vulnerable groups (Haimson et al., 2021; Thach et al., 2022, 2024; Liang et al., 2021). For example, Blackwell et al. examined HeartMob users’ experiences on a platform built for those most affected by severe online abuse  (Blackwell et al., 2017). They argue that platforms’ design and moderation must integrate vulnerable users’ unique needs to “fully” address online harassment. This is also the rationale for our emphasis on examining the diverse experiences of users from marginalized identities with flagging mechanisms. We explore participants’ concerns about flagging with a particular focus on fears of retaliation, concerns about privacy, and misuse of flagging mechanisms.
When considering the design of moderation systems, it is also important to consider end-users’ preferences for control and transparency. Currently, most social media platforms do not offer adequate visibility into the rules and the decision-making processes behind their moderation decisions  (Suzor et al., 2019). Sometimes, transparency challenges are a result of issues with existing moderation interfaces that, for example, do not allow for moderator transparency  (Juneja et al., 2020) or transparency about what decisions are made by automation or AI versus human moderators  (Ozanne et al., 2022). Some researchers have examined how different configurations of moderation tool design shape the control that users perceive over their social media feeds (Jhaver et al., 2022, 2023). Additionally, when moderation systems indicate the reasoning behind content regulation decisions, user attitude toward decisions may improve, resulting in higher-quality user posts  (Jhaver et al., 2019c, a; Kim, 2009) or generally increased trust  (Brunk et al., 2019). Flagging systems similarly lack transparency as platforms are often unclear about when they act on flagged content and what actions they perform. Such practices result in user perspectives being overlooked and unheard. For this reason, we offer empirical insights about how users view individual agency in flagging processes and examine their transparency needs.
2.3. Flagging Mechanisms on Social Media Platforms
While it is unclear when flags were first implemented on social media platforms, the notion of raising a concern to a community leader or administrator has long been used as a strategy for expressing dissatisfaction civilly. Flags are a crucial means for users to express such dissatisfaction in online environments.
Previous research examined flagging from end-users’ perspectives as creators of content, especially in the context of how users develop “folk theories” (DeVito et al., 2017) about moderation decisions when platforms fail to provide adequate or any explanations for those decisions. The opaqueness of flagging has led to speculations about platforms making controversial decisions about flagged content without the flagged user’s knowledge, such as with shadowbanning (where platforms hide a user’s content from others’ feeds without removing it entirely)  (West, 2018). This has raised concerns about the use of flagging to silence (Are, 2023) or otherwise censor users (Peterson et al., 2013). Our study expands on this work by specifically examining the perspectives of users when they are flagging others’ content rather than being flagged themselves.
The definition of how “flaggable” a post is, or its “flaggability,” for users may differ from platforms’ definitions of what is inappropriate. For example, users on gaming platforms utilized flags to report what they defined as toxic content but also flagged teammates to whom they attributed game losses (Kou and Gui, 2021). What’s more, what is interpreted as harmful enough to flag may differ from user to user (Jhaver et al., 2018a), and platforms have differing guidelines on what is deemed inappropriate and thus flaggable (Pater et al., 2016). These findings highlight a gap in our understanding of how users navigate conflicts between their values and platform guidelines for flagging and whether such appropriations of flags occur on social media platforms. Thus, we examine the decision-making process behind how users choose to flag inappropriate content online. Drawing from prior literature, we refer to inappropriate content as a multitude of behaviors (Scheuerman et al., 2021) that include but are not limited to trolling  (Gerlitz and Helmond, 2013; Phillips, 2015), invasion of privacy  (Sambasivan et al., 2019; Chen et al., 2023; Boyd, 2008; Aïmeur and Schőnfeld, 2011), public shaming  (Vaccaro et al., 2021; Ronson, 2015; Goldman, 2015), and interpersonal harm (Riedl et al., 2021).
At the moment, most major social media platforms provide users with the option to flag content. However, the actions users must take after hitting the initial button to flag content vary. For instance, when flagging content on YouTube, a user just clicks on the “report” button, selects a category of harm, and then is thanked for their efforts. At this point, users receive no notifications about the outcome of flagging such content. The flagged content may also continue to appear in their search results. Other platforms such as TikTok, X (previously Twitter), and Instagram offer the option for users to block or mute (i.e., not recommend the content of) accounts they flag or that are responsible for posting content they flagged. Facebook, at the moment, has the most objectively visible post-flagging process, providing users with a specific portal to check the status of their flags (Fig. 2). While previous studies have examined interventions similar to flags to mitigate online harm  (Katsaros et al., 2022; Chang and Danescu-Niculescu-Mizil, 2019; Jhaver et al., 2023), we expand this work in our focus on all three stages of the flagging process: 1) before 2) during and 3) after submitting a flag.
3. Methods
3.1. Participant Recruitment
Our university’s IRB [^1] We will specify the University name after the anonymous peer review is completed. approved this study on Jan 31, 2023. We recruited participants by posting on Twitter, NextDoor, and Craigslist. In selecting these platforms to place our recruitment calls on, we aimed to reach out to a diverse set of potential participants. We asked candidates to submit an online form that included questions about whether they use social media daily, which social media sites they use, whether they have encountered toxic content on social media, whether they flagged a post in the past week and their demographic information. We included an option to attach a screenshot for a post the users recently flagged. We also included an open-ended question: “What is your perspective on how social media platforms can improve flagging mechanisms?”
We reviewed submissions and screened for responses with legitimate and thoughtful answers. In selecting interview candidates, we drew inspiration from prior literature that advocates for researchers and practitioners to give voice to the perspectives of users who are more likely to be vulnerable to online harm  (Duggan, 2017; Musgrave et al., 2022; Blackwell et al., 2017) as well as literature showing such users are more likely to be exposed to content that is harmful (Duggan, 2017; Jhaver et al., 2018b). Therefore, we prioritized offering interviews to candidates who self-identified as belonging to marginalized groups, such as Black and LGBTQ+ users. Table 2 shows the demographic details of our participants.
All participants indicated that they had flagged at least one social media post within the past month of their interview date. We used this experience of flagging to ground our interview discussions and let participants provide more specific feedback to our queries about their understanding of the purpose and operations of flags, their motivations for flagging, and their concerns with flagging.
3.2. Data Collection and Analysis
In total, we collected data through 25 semi-structured interviews with social media users. Participants were compensated with \$20 USD for their time. Interviews lasted between 30 and 90 minutes and were all conducted remotely over Zoom. All interviews were recorded and transcribed. At the start of each interview, we asked participants for explicit verbal consent for their audio and/or video to be recorded for us to conduct analysis. Table 2 (Appendix A) shows the demographic details of our participants.
While participants were only recruited if they self-reported that they flagged content, we provided a common definition of flagging to ensure the research team and participants were on the same page. We defined flagging as “when someone reports a post or accounts for toxic content.” We intentionally left this definition flexible to ensure candidates understood we were speaking about flagging experiences in the general context of reporting online harm. During our interviews, we asked participants about the platforms they used and why they used them to understand the contexts in which they perform flagging. Next, we asked them to describe at least one instance in which they flagged content and inquired about their understanding of how flagging mechanisms operate. We also asked questions to elicit thoughts about the division of responsibility for managing inappropriate posts between platforms and users. We prompted participants with questions aimed at stimulating wider reflection about the flagging design space, often by provoking them with hypothetical changes to flagging designs. While we did not focus on a specific platform during recruiting, the three most common platforms participants discussed were Facebook, Instagram, and Twitter (now known as X). [^2] We refer to the platform X as Twitter because Twitter had not yet made the name change when we began recruiting participants. Specific details about which platforms each participant frequented can be found in Table 2 (Appendix A).
We read and uploaded interview transcripts to Nvivo, [^3] https://lumivero.com/ a cross-platform app for qualitative research for analysis. We applied interpretive qualitative analysis to all interview transcripts  (Merriam, 2002), reading and coding each interview soon after it was conducted. We “open-coded”  (Charmaz, 2006) interviews in multiple iterative rounds, beginning on a line-by-line basis with codes sticking close to the transcript data. This first round of coding generated codes such as “wanting to understand platforms’ timeline for taking action with flags” and “platforms should prioritize by quantity of flags.” In subsequent rounds, we merged related codes to generate higher-level themes (e.g., a need for platform transparency about flagging processes and the necessity of flags) and identified connections between themes. Throughout coding, we also engaged in regular discussions and memo writing, which helped us achieve deeper reflections and stay alert to emerging themes. After we finished processing twenty interviews, we reassessed our data and conducted five more interviews to better flesh out themes such as “value of flagging” and “responsibility for flagging.” At that point, our analysis reached theoretical saturation, and we concluded our data collection.
4. Findings
In this section, we present findings on users’ mental models of flagging (RQ1), motivations to flag (RQ2), and cognitive and security concerns (RQ3). We begin by exploring how users learned about flagging and their understanding of its mechanisms. Next, we discuss key motivations for flagging, such as preventing harm, social pressure, and civic duty, as well as factors that discourage further flagging. We also present participants’ feedback on flag design innovations, like making flags public or granting privileges to certain users. Finally, we address concerns about the burdens of flagging and potential security risks.
4.1. RQ1: What are users’ mental models of how platforms implement flagging mechanisms, and how do they develop these models?
In the following section, we detail findings about how our participants learned about the existence of flagging and developed an understanding of how it functions.
4.1.1. Introduction to flagging
We found that participants first learned about flagging in social settings (e.g., when with friends) or while alone (i.e., exploring site functions while on a social media platform). Participants sometimes began flagging after watching friends flag (P9, P11, and P17). Others were introduced in differing social settings: P8 and P15 were taught about flagging in school, and P15 was encouraged to flag racist content by his history teacher as a way of taking action: “I went to talk to her about [the racist content] and then she… explained that if someone did this to you…you could take action apart from just talking and nobody listening to you, you can actually flag.” Such anecdotes show how information about flagging may be spread via real-life relationships users have.
Several other participants indicated that they learned about flagging while exploring platform options for addressing harm. These participants expressed a common sentiment that on most platforms, the user interface for flagging mechanisms made it easy for them to figure out what the function was for:
“The UI is so well designed that the user finds the report button very easily and very quickly as soon as they join the platform, and it’s also good because every platform now is kind of expected to have that functionality. So you look for it when you want to report something.” (P12)
Participants also heard about flagging as an option through social media posts and news articles. For instance, they learned through posts directly on the platform (P2, P7, and P14) or, more specifically, from creators who mentioned flagging while discussing negative experiences with platforms moderating their content (P4 and P12). While these were the most common ways for our participants to hear about flagging, platform policies could also make users aware of the choice they have to flag. However, only P3 revealed that she personally read platform policies to learn about flagging.
4.1.2. Deciding what to flag
While participants shared a common understanding of the main steps involved in flagging content, they had varied interpretations regarding the specifics of some steps. Participants described flagging as involving three main steps: (1) before flagging (i.e., deciding if content should be flagged), (2) during flagging (e.g., categorizing the content and explaining why they are flagging), and (3) after flagging (i.e., receiving notifications about the outcome of their flag) (see Figure 3). P23 describes this process:
“You stumble upon a post that you deem violates the community standards, from there platforms they have a dropdown menu… [then it] shows the different reasons why you would report such content. And then once you submit that, usually, you’d get a notification saying ‘Thank you for reporting this. We’ll let you know if this violates our community guidelines,’ and then you wait and see.”’
Participants flagged a diverse set of posts that they viewed as inappropriate, following their judgment or per platform policies. Awareness of platform policies primarily came from reading the options presented when first flagging, as few of our participants went out of their way to read platform community guidelines or other policies for similar purposes (P3, P15). There were common categories of content that participants described as inappropriate, such as sexual harassment (P16, P18, P21), domestic abuse (P19), impersonation (P24), spam (P12, P16, P20), nudity (P20), racism (P2, P3, P15), and sexism (P1 and P25). However, participants could not easily determine whether all types of content were flaggable due to ambiguities in platform policies. For instance, P3 pointed out:
“When someone’s saying something that’s not true [such as] victims are lying when it comes to sexual assault… I wish that could be reported, but I feel like even pre-Elon Musk Twitter would just see that as an expression of opinion.” (P23)
Participants spoke about several decision-making processes they applied in these cases. While some opted to flag content that could be inappropriate, others took more of a minimalistic approach. For instance, P9 flagged posts every time she saw something she did not like, while P10 would hold himself back because “[he didn’t] have enough reasons to do that.” In the sexual assault case mentioned above, P23 stated she would opt to not flag such content because she perceived it would not be taken down since it could not be categorized as what the platform deemed inappropriate.
Once participants determined that the content was inappropriate, they would initiate a flag submission by clicking the flag button for that content, which was typically followed by selecting the harmful category (and subcategories, on some platforms) they deemed relevant.
4.1.3. Outcomes of flagging
After participants submitted their flags, they described a range of outcomes. In successful cases, some participants reported being notified that the platform agreed with their flags and that the flagged content was removed (P4, P6, and P7). However, several participants stated that not all platforms sent notifications about the outcomes of flags. For instance, P12 said that “the comment [she flagged] just went away for [her], it just disappeared, and YouTube never followed up with [her] on it.” Additionally, among platforms that did provide notifications, participants reported that platform practices have been inconsistent over time: P2 and P22 never heard back from the platform admins when flagging on Facebook, while P23, whom we recruited at a later date, received outcome notifications. Participants also stated frustration with where notifications were placed, as platforms could notify participants as part of their typical notification feed or in a separate portal specific to flagging. Per P23:
“I wish it wasn’t through a separate portal… I wish you could… put out this feature where they would give you an actual notification. You can click on the notification, and it will take you to like the update.”
4.1.4. Performing other moderation actions while flagging
After flagging content, our participants often described engaging in additional moderation measures, varying from blocking an account to messaging the account posting inappropriate content. For instance, P16 always direct messages accounts that post harmful content and only flags the content if he is unsuccessful in convincing the account owner to take it down: “What I wanted to do first was inform the person nicely. Well, since I didn’t get a reply, I don’t really like flagging but it’s then time I see the effect of the flag.”
P16 claimed that if he saw many comments denouncing a post, he would refrain from flagging because he felt it would be redundant, assuming that others would have flagged that post. P15 and a friend once personally confronted members of an organization behind an account posting harmful content, successfully eliciting an apology and watching them remove the content; as a caveat, P15 described their identity as an activist, which likely attributed to such a proactive strategy regarding flagging. P19 ‘likes’ offensive posts because he believes they will go viral, encouraging others to flag them. In this case, P19 exploited the ‘like functionality’ to trigger moderation.
4.1.5. Secrecy in the flagging process
While participants could describe the flagging process from the user end in detail, they had many questions about how it worked on the side of platforms. In fact, they seemed to relish the opportunity to ask questions when presented with the chance to learn more about flagging from this perspective because certain aspects of the flagging process were akin to a black box for them.
Participants were curious about how “penalties” or consequences for those posts or accounts taken down were determined. Questions about consequences ranged from P15, who wondered what actions were taken before platforms send out automated messages to those who flag content, to P10, who wondered whether the penalties related to content categories or if there was any appeal process for those whose content is flagged.
P20 wanted to know why it seemed that platforms do not have a consistent way of notifying both the account whose content is flagged and the person flagging it about the results, theorizing from experience that platforms may prioritize notifying the account that is flagged. We now know this theory is likely not true across all social media platforms due to shadowbanning and other instances where users who post content have been unfairly censored or even banned without notification  (West, 2018; Are, 2023). However, P20’s inquiries highlight important questions about why this type of secrecy occurs. P20 was also curious about how platforms navigate ambiguity within different types of harmful content (e.g., the difference between teasing and outright harassment) when making flag review decisions.
Ultimately, the predominant question among all participants concerned the decision-making process. Participants asked variations of the question, “How do platforms decide what flagged content is taken down vs kept?” Their curiosity ranged from P22’s “What is going on behind the scenes?” to P21’s query about “what the first people on the ground see.”
When encouraged to suggest solutions for enacting greater transparency regarding flagging, participants’ responses alluded to a system similar to package tracking, where they would be notified immediately after they flag content that the flag submission was successful, informed about the current stage of the flag review, and then be updated about the final decision. As P3 shared:
“Tell me at what point there’s automated decision-making. Tell me what point there are humans. Tell me what the overall policy is. Tell me the percentage. Give me data about how many things were reported.”
The other pressing question on participants’ minds concerned why it seemed to take so long for platforms to act after they flagged content. Seven participants asked this question. In describing their frustration, they recognized that platforms are faced with a large volume of content to moderate but did not consider that as an acceptable reason for weeks-long delays or no responses to their flagged content. P20 voiced this concern constructively: “I would ask them how long should it take?…Should we expect a message from them?”
4.1.6. Speculating about platform motivations.
Dissatisfying experiences lead some participants to believe that flagging rarely leads to removing flagged content and to speculate about platforms’ motivations in providing users with the option to flag at all. Participants surmised that platforms let users flag content for their own benefit and to protect users. P4 shared that platforms have so much content to moderate that they have little choice but to outsource the work to their users: “There is so much data there that’s being posted, that it’s impossible to have like a team that moderates everything.” Similarly, P10 speculated that platforms benefit from users flagging content: they collect flagging data and use it as feedback to improve content recommendations or platform-wide moderation by funneling information to the correct decision-maker. P1 and P3 suspected that legal reasons and pressure from sources such as the Congress drive platforms to offer flags. Per P1: “I think it’s for legal and compliance reasons.”
Participants expressed varying degrees of trust in whether flagging as a feature is designed to truly safeguard end-users. At one end, P20 thought that platforms offered flags as a smokescreen:
“I feel like they want us to feel like we can regulate what you see. They want to make us feel like we have an option in what we see and what we interact with. We have a say in the company.” - P20
Along this line, P16 speculated that platforms may be motivated to ignore flags that could benefit their business models, such as not taking offensive content down because keeping it up may help “generate a lot of user traffic.”
Conversely, some participants reasoned that platforms offer flags to genuinely protect and empower users. They felt that platforms are motivated to protect and include users to keep the space they created safe (P1, P6, P12, and P18).
4.2. RQ2: What motivates users to engage in flagging on social media sites?
Our participants were motivated to flag content for various reasons, the most common being to protect themselves and others from exposure to inappropriate posts. Participants’ moral values guided them in deciding whether to flag specific posts. They further asserted that they independently made decisions to flag, even if they were pointed to potentially inappropriate content by others.
4.2.1. Protecting self and others from online harms
All participants flagged content they encountered on their feeds that they felt was harmful. However, they maintained that the decision to flag content was based on their ethical judgments, not what platforms explicitly define as violating their guidelines. Participants were motivated by values centered around the self, family and friends, and their communities. Six of them reported flagging content that considerably violated their personal values; however, their reasoning for doing so varied. For example, P10 and P19 considered their own values when deciding whether the content is “right or wrong” to flag. P19 defined his conscience as the “moral standard” for his decisions to flag: “As long as it’s up to my conscience that this is right or this is not right…I flag because, according to me, going to my judgment, this post is bad.” P14 and P20 explicitly considered first and foremost protecting their friends and family from content they find offensive or burdensome when flagging. P20 explained, “It’s coming to protection. I’m protecting my kids, those people that I know.”
Others were motivated to flag content they believed would harm their community, often citing examples like racism or explicit material and considering the protection of vulnerable groups, e.g., children. P9 argued that flagging in “social media is not about just flagging…because what you’re doing will help someone. One day you don’t know if the person you’re helping is your friend or your brother.”
4.2.2. Submitting flags due to social pressure
Participants also admitted to flagging content when prompted by others. They described getting invited to flag via seeing others’ comments, through direct messages (DMs), or hearing from friends through other channels of communication. P1 said, “So my friend had messaged us in the group thread and was like, hey, this is really bothering me. Can you all please report and join me?”
Similarly, P6, P15, and P16 were contacted directly by people they knew who asked them to flag something. P12 noted that though she flagged comments because others urged her to, she did so only if she agreed with others’ rationale. P19, however, was less strict when deciding whether to flag, claiming, “I might flag the post without even checking it” when someone asks him to flag something. Thus, participants showed varying degrees of compliance to social pressure in their flagging behaviors. Some participants were motivated to flag content by others who shared their identity characteristics. For example, P5 said:
“I have been encouraged particularly by other natives and other queer people and trans people to go ahead and report something. Usually, it’s a prominent figure who said something that really crosses a big line.”
All participants who were prompted to look into harmful content and flag it by others agreed to examine the content, but most emphasized that the final decision to flag the content was theirs alone. The only exceptions were cases where participants flagged to appease their friends who expressed their passion for sanctioning the post in question. In such cases, P19 and P21 said they would flag the content because they trusted their friends. Per P19: “Because I knew she said it’s bad, so I will actually flag it.”
4.2.3. Perceiving flagging as an obligation
While all participants believed platforms are responsible for proactively detecting and regulating offensive content (which many erroneously referred to as platforms’ “flagging”), most also thought individual users also have a responsibility to flag content.
More than half of our participants believed that every user is responsible for flagging content. This belief was strongly tied to their reasoning to flag content to protect others. P8 expressed: “I think that it is a collective duty to keep social media space clean, and so it’s everyone’s responsibility to flag all inappropriate things.” P13 and P22 shared P8’s sentiments and used this cleaning analogy to argue that everyone is responsible for contributing:
“If everyone does it, then the world would be a better place…If there are some plastic bottles in the street and someone picks them up and puts them in the recycling — it is not going to remove all the plastic from the environment, save the turtles, whatever. But it’s a start, and everyone should do their part no matter how small it is.” (P22)
4.2.4. Flagging as an avenue for user expression within content moderation systems
Participants valued flagging as a mechanism to offer their feedback. In fact, several expressed that flagging was a “right” they had as users on social media platforms (P2, P10, P19, and P20). Despite not knowing the outcome of their flags or finding that flagged content was not removed, participants remained optimistic and confident about the value of flags in expressing their voices.
“Maybe it’s not really helpful, or maybe it’s merely making me feel better. But if something is done about the fake posts, which sometimes happened, at least in terms of Facebook, then I’m probably helping others avoid potentially very dangerous information.” (P3)
Similarly, other participants believed in the value of flagging as a way to interact with the moderation system. P12 thought that it offered “an excellent way to moderate the space that a lot of people use every day” and found the value in flags even greater because of the alternative, i.e., having no option to give input at all. In her words, “thinking of a world where there’s no reporting, gosh…there must have been a lot of toxic language thrown around.”
4.2.5. Issues with inaction and communication
While platforms’ inaction dominated users’ sensemaking around flagging outcomes (sec. 4.1.3), it also demotivated users from engaging in flags. For some participants, experiences flagging posts were particularly disappointing. P2 received an automated message from a bot immediately after flagging content that acknowledged receipt of the flag, while P15 and P20 also received some level of communication from the platforms on which they flagged content. However, none of the posts these participants flagged were taken down even weeks after they flagged them. Participants’ sentiments rang strongly of frustration and feeling wronged. P2 and P15 used almost the exact same words about platforms communicating with empty promises. P2 said:
“They told me that they will take action and that they will remove both, but they haven’t done it up to now…It’s about a week, and they haven’t done anything. They should have done something a long time ago, like on the same day [it was flagged].”
P22 flagged both an inappropriate post and the account that posted it. Though he received a notification that the post he flagged was taken down, he received no subsequent communication about the account he reported. Therefore, in this case, P22 considered the communication as lacking.
Several participants received no communication from platforms after flagging content. Of this group, the flagged content seemed to disappear for P6, P12, and P17. However, none of these participants followed up with the content, claiming they noticed only that it stopped showing up in their feed. P12 reasons that the lack of follow-up from platforms could be due to the high volume of reports.
“The comment just went away for me like it just disappeared, and YouTube never followed up with me on it. I can’t imagine they would because if I am anything to go by, they must get hundreds and hundreds of reports.”
Without any feedback from platforms, participants created their own folk theories about the effects of flagging. For example, P4 suspected that Twitter automatically blocks accounts of flagged content, preventing users from seeing any content from that account again in their feed. Therefore, P4 felt blindsided when unable to see the results of her flagging efforts.
“I don’t know if it’s working well because I actually don’t know if it’s working. I have no idea what happened to that account…I have no means to actually see the status of my reporting.”
While some participants could shake off their frustration as they continued to use platforms, the experiences of flagging content discouraged others to a greater extent. These participants stated that negative experiences with having their flags ignored and a lack of communication from platforms decreased their motivation and made them less likely to flag content in the future. P1 developed a strategy when flagging to avoid wasting their effort: “I have to pick and choose because I find that a lot of times when I flag, things get ignored.” P2, P3, and P9 developed a more grim, fatalistic take on flagging, expecting platforms to ignore most flags. In sum, prompt and clear feedback is essential for users to feel heard and be motivated to continue flagging efforts.
4.3. RQ3: What are users’ concerns about the cognitive load and privacy risks of flagging content?
Participants cited many factors that led to their negative perceptions of flagging mechanisms and discouraged them from flagging in the future. Notably, a large majority of participants had dissatisfying experiences when flagging content, feeling burdened by the extra steps they believed flagging involved and frustrated with platforms’ lack of transparency.
4.3.1. Cognitive burden of submitting flags
Participants were distinctly aware of how flagging as a mechanism places the burden of moderation on them. Those with underrepresented identities in our sample especially emphasized the additional burden of flagging they face. For example, P1 observed that they end up seeing more harmful content due to their identity characteristics: “I will flag things because, unfortunately, if you’re a queer person, if you’re a trans person, if you’re a person of color, if you’re a woman, you will regularly encounter really heinous comments.”
Several participants noted that flagging takes more effort than it should. As described in Section 4.1.2, categorizing harmful content can be confusing when platforms’ definitions of inappropriate content categories are unclear and fail to cover all types of norm violations. This burden is exacerbated by platforms’ differing flagging policies, resulting in different categorizations. As P25 proposes:
“I wish you could just click a button, report it, and it would just go to whatever algorithm or software that determines this does or doesn’t meet our requirements [and] maybe you could have the option to comment specifically.”
4.3.2. Platforms’ flagging responsibilities
Some participants believed that the burdens of flagging, as described in the previous section, are exacerbated by platforms not carrying out their flagging responsibilities. These participants agreed that platforms have a responsibility to make flagging mechanisms accessible and to take action on inappropriate content that is flagged. According to P12, “Everybody has the responsibility [to flag], but the responsibility of taking action towards it is of the platforms.” In this way, platforms’ responsibilities seem analogous to a restaurant owner’s responsibility to take action to protect their business from rowdy customers who disturb others.
However, there is nuance in the extent to which that responsibility applies and what it entails. For instance, P12 just wanted platforms to “follow their own policy” and called for platforms to be transparent when following their own policies for flagged content, indicating that it did not matter to her how they conducted flag review (e.g., outsourcing to third parties is acceptable). Similarly, P3 expected that platforms would take at least some action on flagged content that is egregiously offensive to the point where it “would not really be up for debate.” However, it is notable that P3 explained they only came to this conclusion due to having low expectations of platforms taking action against inappropriate content. P15 believed that platforms’ responsibility extends further, i.e., platforms like Facebook should assume legal responsibility for hosting harmful content such as death threats.
Participants also considered that platforms have the capacity and resources to review most posted content themselves rather than relying on end-users. They noted that platforms already deploy human moderators and automated software for content review, and this infrastructure should be leveraged further. For example, P15 felt that identifying inappropriate content is “not that much of an effort in the first place” for platforms since it is “what they’re supposed to do.”
4.3.3. Prioritization Order of Flag Reviews
Prior research shows that Black, Indigenous, and people of color (BIPOC), LBGTQIA+, disabled, plus-sized individuals and sex workers experience more frequent and severe moderation consequences than other users  (Haimson et al., 2021; Ma and Kou, 2022; Are, 2023; Lyu et al., 2024). Given these inequities in moderation outcomes, we explored our participants’ attitudes toward policies that dictate the prioritization order of flag reviews.
First, we prompted participants with the question: should users be considered equally when flagging? When presented with the idea of a “super-flagger status,” i.e., individuals given higher priority for review when they flag content, most participants said that all users should not be considered equally, though not all agreed on the details of that prioritization. P10, P15, P16, and P18 argued that platforms should prioritize reviewing the posts flagged by influential figures such as celebrities and political leaders because their popularity suggests that many users would trust their flagging actions. P15 said:
“If they flag something right now, it should be taken serious much quicker. [The likelihood] is very high because of that large audience they reach. Basically, they have a large audience, so they should be given priority.”
Other participants asserted that users should be prioritized based on their reporting history (P4, P5, P12, and P19). For instance, if users abuse flagging privileges, the content they flag in the future should be lower in priority for the platform to review. P19 also argued for prioritizing users’ flags based on the amount of time they spend on the platform.
In contrast, some participants argued that all flag contributors should be considered equally. They offered three beliefs to support this argument: the ideal of equal rights for all suggests that all social media flaggers also be treated fundamentally equally; having more followers does not qualify one to have more privileges when flagging; and prioritizing by user as a strategy is not effective. P20 considered the situation from a business perspective: “I feel like we are all promoting Instagram by using it. So we are all consumers, and we are all adding revenue to the company. So they should consider us equally.”
We further explored this topic with our participants by inquiring: should types of flags be considered equally? Our participants’ responses to this question show a lack of consensus on how flag reviews with different categories of rule violations should be prioritized. Five participants maintained that not all categories of flagged content should be considered equally (P1, P9, P15, P16, and P17). P1 and P9 mentioned harmful content that is potentially life-threatening to be reviewed with the highest priority, while P15 thought racial discrimination should be at the top of the list.
Many participants opposed this view and instead held that all flags should be considered equally (P2, P8, P14, P17, P18, P19, and P20). P18 had a simple yet strong reason that flags should be considered equally: “If something is not good for the community, it’s not good for the community.” Similarly, P8 viewed flagging itself as a simple mechanism to generate candidates for the moderator queue, saying that “flagging is just to bring attention to review it.” Some participants felt that establishing a priority order for reviewing different content categories is untenable. For example, P20 could not determine how the flagging system could prioritize handling racism vs sexism because of her firsthand experiences of being discriminated against both as a woman and a Black person.
4.3.4. Privacy for personally identifiable information
In addition to concerns about the cognitive burdens of flagging, participants also surfaced several privacy-related concerns. Participants expressed apprehensions about where flagged submissions go for processing and retaliation from opposing parties. Several participants did not want their personally identifiable information to be revealed when they flagged content. P21 expressed concern about human moderators or staff at platforms seeing their information but comforted herself with the reasoning that “these people are way too busy to be counter-checking every account.” P5 also worried about the trustworthiness of moderators behind the scenes and speculated that those with ill intentions could get an “entry-level position” and “pass the background check” to become a moderator.
In contrast, both P18 and P22 had few to no concerns about their personal information being leaked when they flagged content. P18 preferred that the platform keep information about past content he flagged so that similar content would not show up in his feed in the future.
4.3.5. Retaliation from flagged users
Although in the minority, some participants also related negative experiences where owners of accounts or content they flagged retaliated, causing them to feel unsafe. P20 was concerned that users from marginalized groups who get flagged would assume that flaggers report them because they are discriminating against their identities even when she flags due to other reasons.
“For example, reporting a page, I feel like they might think we are biased toward that certain type of group, maybe they’re not straight, and we’re not supporting them. Maybe we’re homophobic.”
P12 developed a habit of blocking accounts after receiving hateful comments when speaking out against a post and then flagging its content: “They (account owner of flagged content) messaged me privately, and they called me names. So I ended up having to block and then report them.”
P12’s anecdote raises an important consideration. Though flagging is anonymous at the moment, in the aftermath, an array of concerns with retaliation arise that must be examined before new features are implemented, e.g., showing one’s followers the posts or accounts one flagged. These concerns may be exacerbated if flagging were to become more transparent. To further explore this topic, we prompted our participants with the question of: who should see flags?.
Given the hypothetical option of changing who can see when they flag content, several participants decided that all other users should be able to see when they flag content. In return, they also expected to see other users’ flags. The prevalent sentiment for implementing this change was the belief many participants had that the more times an inappropriate post is flagged, the more likely it is to be taken down. Per P4: “With only one flag in…it means maybe moderators will not pay attention to it, but with other people flagging it means actually some action will be taken.” As such, P7 added that public flags would make platforms more accountable for promptly reviewing the flagged posts. He claimed, “If a post is being flagged, I think everybody has the right to know.”
Other participants suggested alternative approaches, such as having only a subset of people be able to see the content they flag (P9, P3, P19, and P21). P3 preferred that only a few of her followers, such as her close friends, could see the content she flags to avoid conflict since she is followed by people she has “very different views from.” P19 supported having his followers see the content he flags but wanted to be able to toggle this feature on and off in settings. Like P19, P21 preferred that people saw only some types of content she flags since she could be “petty” when flagging some content but passionate at other times about flagging harmful content, depending on the topic.
4.3.6. Misuse of Flags
Prompting participants about changes in flag designs also brought about concerns regarding the misuse of flags. Participants were concerned about the misuse of visible flags. P17 speculated that some might flag content to gain greater visibility on the site, assuming their activity would appear in their followers’ feeds. Participants also held concerns about the misuse of flags in their current anonymous nature. P5 noted experiences where users of marginalized identities who speak out had their accounts flagged and subsequently suspended. From these anecdotes, she became worried that flagging could be an avenue for exacerbating biases against marginalized people:
“I’ve had friends and colleagues who have been targeted by hateful people who all report the poster’s account for spam or something else fake, and then that person gets their account suspended…so it can be weaponized against marginalized people.” (P5)
P3 also recounted problematic past occurrences where groups of “trolls” on platforms such as Reddit banded together to suspend accounts they felt did not ‘belong’ in a certain online community.
5. Discussion
5.1. The Need For and Importance of Flags
Our analysis suggests that users interact with flagging mechanisms in complex ways, and ideas about how they can be improved are plentiful and diverse. While outcomes of flagging are often perceived as disappointing by users such as our participants, our analysis uncovered a strong user sentiment that flagging mechanisms are vital to their experience on social media platforms. Our participants described flagging as one of the core avenues for users to express dissatisfaction with content they deemed inappropriate on platforms. It is necessary because it provides a unique function for these users to anonymously report content directly to platforms to have it removed from the entire platform. Besides flagging, our participants also noted using personal moderation tools, such as blocking mechanisms, to protect themselves from exposure to inappropriate content. Our analysis highlights a distinction that users make among flagging mechanisms and other personal moderation tools like blocking: participants often flag to protect others from exposure to inappropriate posts, but they may primarily deploy blocking to protect themselves (Jhaver et al., 2023).
Thus, flagging is not an isolated tool; future research should examine in greater depth how users deploy it alongside or in place of other moderation mechanisms and actions. Such work may benefit from our conceptualization of the three stages of flagging shown in Figure 3 as other mechanisms would present different affordances, procedural insights, cognitive burdens, and privacy considerations in their analogous temporal stages. Further, certain moderation mechanisms are designed for the integration of specific stages of the process (e.g., blocking a user after flagging). In this way, we can also see how flagging fits into the ecosystem of moderation mechanisms, such as blocking or muting accounts and even more personalized moderation mechanisms that only remove content from a user’s individual feed rather than the platform itself.
Figure 3. Visualizing three stages of flagging.
The anonymity built into flagging mechanisms allows users to avoid retaliation by not having to confront perpetrators who post inappropriate content. In some cases where our participants attempted to avoid flagging by directly commenting on an inappropriate post or directly messaging a user (i.e., requesting the outright removal of another user’s content), they were met with inappropriate comments or no response at all. In these cases, they needed to resort to flagging as it was the only other option they had to request for the content to be removed. This extends our prior understanding of how users leverage their online social connections and content moderation tools (Wright, 2022; Jhaver et al., 2023) to address online harms  (Scheuerman et al., 2021, 2018).
Flagging is a critical part of users’ social media experiences because it is embedded in many social interactions users have on and offline. For example, our participants told stories of learning about the function of flags through word-of-mouth and seeing others flag content. These instances demonstrate that the influences and efforts that shape users’ decisions to flag content stretch beyond interactions on specific social media platforms. What’s more, our findings about how our participants were often encouraged to flag in an effort to protect others or to “team up” by mass-flagging inappropriate content further demonstrates how flagging is actively employed by users for critical social functions such as protecting their friends or reinforcing commonly shared values about what is inappropriate. These insights add to ongoing conversations about how social media users leverage their critical postdigital literacies to create new cultural and ideological meanings of platform-implemented features like flagging  (Hagh, 2024) and how users interact with platform governance  (Chan Kim and Mauborgne, 1998; Ostrom, 1990). In sum, flags serve as a mechanism deeply embedded into users’ social media interactions, providing them a crucial opportunity to report inappropriate content anonymously.
5.2. Platform Versus User Responsibilities for Regulating Inappropriate Content
Our analysis suggests that users perceive flagging in varied ways: as a right, an obligation, an exploitation, and a commons problem. Many users viewed flagging as a right they have that platforms should uphold. The sentiment of generalized reciprocity, along with users’ inclination to take on nurturing and supportive roles in their communities (Seering et al., 2022), leads them to view flagging as an obligation. However, unlike the metaphorical labels of “custodian” or “janitor” that apply to content moderators (Gillespie, 2018; Seering et al., 2022), regular users’ flagging is more metaphorically aligned with a civilian picking up trash on the street. They may feel obligated to do so to keep their community clean and may even be encouraged to by those around them, but it is ultimately their choice to do so. In fact, the weight of the responsibilities users feel to keep others safe through flagging is significant (Roberts, 2014; Musgrave et al., 2022), and it is further exacerbated when flagging systems fail to address the cognitive load involved in submitting flags. The labor of flagging is not to be overlooked and may easily be exploited. Like previous examples of voluntary human labor of moderators  (Dosono and Semaan, 2019; Matias, 2019; Wohn, 2019; Kiel Long et al., 2017), we confirm through our results that users experience similar feelings of cognitive burden when flagging. Platforms then have a responsibility to ensure that flags are not only offered to users but designed to reduce user burdens. For instance, flag designers could reduce the number of mandatory steps involved in flag submissions. Further, when designing communications regarding flagging, platforms should clarify the utility of flags but not impose them in ways that induce non-use due to psychological reactance (Brehm, 1966).
5.3. Incorporating Transparency in Flagging
Prior research has highlighted the importance of defining what transparency means  (Suzor et al., 2019). Additionally, the implementation of transparency or lack thereof has been examined in the context of content removals, e.g., informing users about their post removals and explaining why they occurred (Jhaver et al., 2019a, c; Brunk et al., 2019) and platform moderation practices  (Juneja et al., 2020). Our analysis shows that platform transparency is also crucial in flagging procedures. We urge for transparency that provides users with satisfactory information about each stage of flagging as part of the platform’s responsibilities in offering flags. Applying the knowledge that people tend to behave better when decision-making processes are perceived as fair and just  (Ostrom, 1990), users may be more inclined to flag if they feel that platforms will respect their time and effort in return. Without such improvements in flags, it will be difficult for platforms to continue relying on user flags to identify inappropriate content. For instance, some of our participants concluded that flagging is simply a “smokescreen” with no platform functionality or support, which led them to stop flagging altogether.
A need for transparency in flagging procedures is further heightened by concerns about the use of automation or related technologies (e.g., LLMs) in making content removal decisions. Users are open to the use of automation in this context, especially after they experience little communication or actions on their flags. However, using such technologies requires greater clarity about when automated approaches are deployed and how human reviewers work alongside automated tools. Our analysis revealed a curiosity about this division of labor at different stages of the flagging pipeline, e.g., who evaluates whether the content has been flagged in the correct reporting category, when humans oversee the decisions made by bots, and how sanctions are determined. More insights into these procedures may help end-users better understand platform constraints, manage their expectations about flag processing at different stages, and reduce flag misuse. Therefore, instead of seamless flagging interfaces, platforms should develop carefully designed seams (Eslami et al., 2015, 2016; Jhaver et al., 2018c) that inform users about how flagging works.
5.4. Design Implications for Improving Flags
Through our study, we surfaced tensions and challenges within all three stages of the flagging process (see Figure 3). We consolidate our analysis into design opportunities organized by each stage of flagging as well as key considerations and examples of feature innovations that may serve as design constraints or otherwise limitations in Table 1. Below, we expand on these design opportunities, highlighting the nuances that should be considered when further designing in these spaces.
# Design Opportunities Considerations Examples of Feature Innovation
Before (determining flaggability)
1 Flagging icons Ensure flags are accessible (i.e., easy to see and select; not embedded within long action menus). Red flag present in the upper right-hand corner of each post.
2 Increasing flagging literacy The purpose of flagging (i.e., to report inappropriate content) should be clearly communicated. A pinned post with clearly defined types of inappropriate content to flag.
3 Conveying information about flagged content Further exploration of innovations to flags that decrease the anonymity of flags should be carefully explored while attending to users’ diverse concerns. Display that “15 other users reported this post” or personalized display of “your friend Sally Wang reported this post.”
During (classifying flagged content)
4 Classification schemes for types of inappropriate content Offer greater clarity about the meanings of included categories; ensure transparency in additions or changes to the classification scheme. Examples of content that fits in each category.
5 Options to provide additional information about flags The design must balance overburdening users to include too much information (e.g., free response) and not providing enough options to indicate nuanced opinions. Optional short-response text boxes or file upload.
After (monitoring flag outcomes)
6 Visualization of flag review procedures This system should be integrated within the platform for accessibility. A package-tracking-like system to showcase what review stage a flag is at.
7 Visualization of platform statistics about flagging To preserve privacy, we may urge platforms to provide aggregate statistics about the types of content that is flagged and subsequently removed or kept. A common dashboard or monthly report sent to users.
Table 1. Design opportunities with corresponding considerations and examples of feature innovations, organized by the three flagging stages — before, during, and after flagging.
5.4.1. Before Flagging
Our analysis shows that users appreciated the ubiquity and similarity of flagging interfaces across different platforms. Their familiarity directed participants where to look for flag icons and submit reports against inappropriate content. Therefore, when designing flagging interfaces, new platforms should leverage this familiarity and deploy flagging icons, placements, and vocabulary in ways that mirror existing platforms. Our findings also suggest that platforms should try to minimize the cognitive labor involved in discovering flag icons and completing flag submissions and reduce the time required to process flags.
Some social media users may be unaware of how to use flags or not fully understand their value in addressing online harm. This lack of guidance can lead users to misuse flagging, such as by using it to simply express that they dislike content or show animosity to a user  (Kou and Gui, 2021). Since we now know that users learn about flagging from diverse sources, such as friends and school teachers, we encourage practitioners to design ways to leverage these community interactions. We encourage the collaboration of platforms with schools, educational leaders, media sources, and content creators to promote awareness and understanding of flags and other safety mechanisms on social media. Additionally, existing platform communications about the purpose of flagging can be improved so that they are more accessible to understand.
Crawford and Gillespie raise the question of whether “some flags are worth more than others,” such as flagged extremist content or content flagged by a selected group of individuals (Crawford and Gillespie, 2016). Our findings about user support for such prioritization are mixed and vary by context. For instance, users described certain cases where flags submitted by active members of a community should be prioritized, as they have a better understanding of what is inappropriate in that setting. In other cases, it was unclear what would qualify content as extreme enough to be prioritized over other types of flagged content. This presents an opportunity to further explore how the status of the person flagging and the type of flagged content should be prioritized, if at all. We recommend practitioners design guidelines in partnership with users and carefully consider how prioritization may mean something different across platforms and communities.
On the other hand, users expected that the popularity of a flag should warrant its prioritization. Users showed interest in modifications that would show the number of people who have previously flagged that content but were concerned about invasion of privacy when probed about further modifications that would show more specific information such as demographics (e.g., 50% of users that flagged this post were Asian) or personal connections (e.g., your friend Bill flagged this post). Some of our participants who engaged in flagging in a public manner (e.g., commenting that a post is inappropriate before escalating to flag) often faced retaliation in the form of threatening messages. Some platforms have already implemented flags that are more public-facing to encourage greater user involvement in identifying misinformation. For instance, on the Chinese social media platform Zihu, users flag content for it to be marked as “controversial” and then have the option to participate in a discussion page to raise questions about the content  (Li, 2022). However, as Li (2022) point out, only users with higher “yan” scores from being more active and having a greater impact on the platform as calculated by Zihu could post in the discussions. We advise that modifications of flags are made with careful consideration of users’ privacy needs, equity expectations, and audience management goals  (Marwick and Boyd, 2011).
5.4.2. During Flagging
In line with prior research (Haimson et al., 2021), our participants expressed concern that individuals from marginalized groups face a greater burden of reporting inappropriate content. At the same time, these groups frequently get falsely reported. Therefore, we recommend that platforms invest in creating accurate ex-ante identification (Grimmelmann, 2015) and regulation of activities that harm these groups. Platforms could partner with these groups to better understand the nuances of online harms (e.g., the use of offensive slang words that outsiders may not recognize (Jhaver et al., 2018b)) and the moderation challenges they face. Additionally, platforms may implement methods for users to flag content that falls outside of the existing flagging rule violation classification schemes.
No aspect of flagging should be overlooked when considering improving the design of flags to reduce the burden of labor placed on users. As our findings confirm, small design decisions, such as requiring users to provide free response justification for flagging a post, can discourage users from continuing to flag content. Thus, we advise that labor-intensive options that require long-form answers or the uploading of information, e.g., screenshots) are optional and implemented with a clear purpose (e.g., encouraging users to provide additional context for flagged content outside the existing classification of inappropriate content.
5.4.3. After Flagging
Our analysis shows that users want to feel that their effort in flagging content is respected. At the moment, users’ lack of ability to track the flagging status meant that even after flagging a post, they have to check the removal status of posts they reported repeatedly. Some platforms promise to notify users when an action will be taken on their flags but fail to do so. Users can become disillusioned with the flagging process and doubt that platforms would take any action in these cases. We recommend that platforms institute information and visualization systems that allow users to easily monitor the status of their flags. Given that most platforms have either no post-flagging dashboards or ones that present rudimentary interfaces other than Facebook (which is also limited in its transparency and hosted separately from the main platform), there is much opportunity for innovation in the design space of building post-flagging mechanisms. In addition to this design opportunity, there is a need for greater transparency in flagging, which factors into existing calls that seek transparency from platforms via policy reform. For instance, scholars have previously called for regulating content moderation  (Langvardt, 2017; Buckley and Schafer, 2022). They also examined how existing policies of the EU Digital Services Act apply to problematic platform behaviors with moderation, such as shadowbanning  (Díaz and Hecht-Felella, 2021; West, 2018) and unjust platform standards for moderation  (Díaz and Hecht-Felella, 2021). Such legal attention should also be applied to flagging reforms.
Finally, we recognize that privacy concerns remain at the forefront of discussions around increasing transparency in flagging mechanisms. Thus, we propose the design opportunity of providing visualizations for aggregated statistics about flagged content.
5.5. Limitations and Future Work
In this study, we conducted semi-structured interviews with 25 social media users who recently engaged in flagging. Though our data helped us examine nuanced user experiences and tensions with flagging in-depth, the qualitative nature of our study design and the size of our sample limited us from generalizing our results and learning about the relative popularity of various design choices in flagging interfaces. Additionally, we recognize that our findings may be limited because online behaviors and perceptions may change over time. Future research may explore user perceptions of different ways of prioritizing flags, establishing controls, requesting cognitive labor, and maintaining user privacy through quantitative and longitudinal methods.
We focused on general end-users’ perspectives in this work in our effort to provide empirical data that contributes to Crawford and Gillespie’s existing conceptualization of flagging  (Crawford and Gillespie, 2016). It would be valuable to examine the dynamics and challenges of flag processing from the perspective of platform administrators and moderators. For example, it might be insightful to learn how false flagging is identified and sanctioned and design solutions that discourage users from submitting flags in bad faith.
We also concur with and acknowledge critical work highlighting the potential misuse of content moderation mechanisms and flagging in particular. Those with marginalized identities are most likely to fall victim to such abuses of flagging  (Musgrave et al., 2022). The misuse may occur by platforms unjustly censoring or hiding flagged content (West, 2018) or result from actions by individual users. It was for this purpose that we intentionally recruited participants with perspectives as members of marginalized communities. The design considerations we provide highlight important concerns when redesigning flags in this context. For instance, while it may be productive to experiment with non-anonymous flags, we must consider what harm public views of flagging information can result in.
Our analysis shows that within the limited vocabulary of flags, users develop decision-making guidelines based on their values. Flagging can be socially motivated for some, who consider or even flag content without deeply examining it because they value the opinion of another person who asked them to flag. This distance between platforms’ and users’ notion of flaggability aligns with prior research on flagging motivations in online gaming communities (Kou and Gui, 2021). Examining this distance systematically could be an exciting avenue for sharpening platforms’ moderation policies and practices to better align with users’ values.
Finally, the labor users provide to platforms in the form of flagging shares similarities with the labor of data labeling  (Gray and Suri, 2019; Morrow et al., 2022) in its voluntary nature and the potential for exposure to psychologically harmful content. Particularly as applications of large language models become more prevalent, we envision possible applications of flagging mechanisms to allow users to report harmful model output (e.g., adding the flag option to ChatGPT).
6. Conclusion
At a time in which user participation in social media moderation is being recognized as increasingly vital (Jhaver et al., 2023, 2022; Vaccaro et al., 2021), our study serves as a starting point for exploring how users report inappropriate content and how mechanisms like flags can be designed to better facilitate these interactions. Our research examined why end-users flag, their experiences with flagging mechanisms, their mental models of the flagging pipeline, and reasons why they may hesitate to flag. We found that users perceive flags not as a stop-gap solution that compensates for platforms’ limited moderation resources but as an essential avenue to voice their disagreements with the curated content. Our participants’ willingness to contribute their time, even going as far as personally contacting the people who post content they flag to take it down, shows users’ investment in addressing online harms. Our analysis shows that flagging serves as an act of generalized reciprocity for users. However, we found a lack of procedural transparency in flagging implementations undermines users’ trust in them. Many questions about how to optimally design flagging interfaces or what role other stakeholders, such as flagged users, moderators, and bots, play remain unanswered. We call for researchers to investigate these questions and examine flags as part of an ecology of communication tools individuals use to address online harm.
Appendix A Participant Sample Description
Table 2 shows the demographic details of our survey participants.
# Age Gender Race Occupation Country Social Media Platforms Used
P1 35 Non-Binary White Digital Content Producer USA Facebook, Instagram, Twitter, Tiktok, Snapchat
P2 32 Female Black Fashion Designer USA Facebook, Instagram, Twitter
P3 53 Female White Professor USA Facebook, Twitter
P4 27 Female White PhD Researcher France Twitter, Whatsapp
P5 33 Non-Binary Native American Editor USA Facebook, Instagram, Twitter, Tumblr
P6 22 Female Black Waitress Canada Facebook
P7 24 Male Black Artist USA Facebook, Instagram
P8 27 Male Black Carpenter USA Facebook, Twitter, Snapchat
P9 22 Male Black Fashion Designer USA Facebook, Twitter
P10 30 Male Black Electrician USA Facebook, Instagram
P11 26 Male Black Plumber USA Instagram, TikTok, Snapchat
P12 20 Female Asian Student Canada Facebook, Instagram, Youtube, Reddit
P13 24 Female Black Student USA Instagram, Twitter
P14 27 Male Black Uber Driver The Netherlands Facebook, Instagram, Twitter
P15 26 Male Black Sales Representative Belgium Facebook, Instagram, Pinterest
P16 35 Male Black Engineer UK Facebook, Instagram, Whatsapp
P17 27 Male Black Accountant UK Facebook, Twitter, Reddit
P18 25 Male Black Sales Agent USA Facebook, Twitter
P19 25 Male Black Manager USA Twitter
P20 34 Female Black Nurse USA Instagram, Twitter, TikTok
P21 26 Female Black Assistant USA Facebook, Instagram, Twitter
P22 25 Male Middle Eastern Recent Graduate USA Facebook, Reddit, Linkedin
P23 22 Female Asian Student USA Instagram, Facebook, Reddit
P24 21 Male Asian Student USA Instagram, Twitter, Reddit
P25 24 Female White Unemployed USA Twitter, Instagram
Table 2. Participants’ Demographic Information